# Retrosynthesis Planning via Worst-path Policy Optimisation in Tree-structured MDPs

### v1

@@ Abstract
Retrosynthesis planning aims to decompose target molecules into available building blocks, forming a synthesis tree where each internal node represents an intermediate compound and each leaf ideally corresponds to a purchasable reactant. However, this tree becomes invalid if any leaf node is not a valid building block, making the planning process vulnerable to the “weakest link” in the synthetic route. Existing methods often optimise for average performance across branches, failing to account for this worst-case sensitivity. In this paper, we reframe retrosynthesis as a worst-path optimisation problem within tree-structured Markov Decision Processes (MDPs). We prove that this formulation admits a unique optimal solution and offers monotonic improvement guarantees. Building on this insight, we introduce Interactive Retrosynthesis Planning (InterRetro), a method that interacts with the tree MDP, learns a value function for worst-path outcomes, and improves its policy through self-imitation, preferentially reinforcing past decisions with high estimated advantage. Empirically, InterRetro achieves state-of-the-art results, solving 100% of targets on the Retro*-190 benchmark, shortening synthetic routes by 4.9%, and achieving promising performance using only 10% of the training data - representing a significant advance in computational retrosynthesis planning.
@@ 1 Introduction
Retrosynthesis aims to identify the reactants needed to synthesise a target molecule with desired properties. As a fundamental task in computer-aided molecular design, retrosynthesis underpins progress in drug discovery and materials science [4, 39]. A key challenge in retrosynthesis is single-step prediction, which involves predicting the reactants for a given product (illustrated in Figure 1). While recent approaches using supervised learning have achieved human-level accuracy in this task [2, 38], their practical applicability is limited, as the suggested reactants are often commercially unavailable. Unlike single-step prediction, multi-step planning recursively decomposes the target into simpler intermediates, aiming to construct a synthetic route whose leaf nodes are purchasable compounds. This process naturally forms a sequential decision-making problem, where early decisions influence future steps and the overall outcome [7, 21].
Most methods tackle this decision-making problem using heuristic search, with Monte Carlo Tree Search (MCTS) being widely adopted for its ability to balance exploration and exploitation [17, 23]. In this framework, a single-step retrosynthesis model suggests candidate reactions and predicts the most likely reactants for a given product. During simulations, MCTS selects reactions by weighing their estimated value against their exploration potential. After simulating a complete synthetic route, the values of the decisions made along the path are updated to guide future searches more effectively. Building on this foundation, recent studies have proposed various enhancements to improve the exploration–exploitation trade-off or to stabilise value estimation [1, 7, 31, 37]. However, these methods heavily rely on time-consuming real-time search, requiring hundreds of model calls for each molecule, limiting their practical utility in large-scale molecular design scenarios.
To improve search efficiency, recent work fine-tunes pre-trained single-step models by imitating decisions made during successful heuristic search trajectories [9, 10, 34]. The key insight is that reaction choices selected after look-ahead planning are often more reliable than those proposed by the original model [22, 28]. By training on these improved decisions, the resulting policies can propose more plausible reactions and accelerate the search for viable synthetic routes. However, this fine-tuning process adapts the model to the distribution of molecules encountered during search, rather than those seen in direct inference. As a result, the model may perform well when used within the search loop but struggle to construct full synthetic routes independently — limiting its utility in settings where fast, search-free planning is desired.
@@
In this paper, we propose a novel perspective: reframing retrosynthesis planning as a worst-path optimisation problem in tree-structured Markov Decision Processes (tree MDPs). We observe that existing approaches typically optimise for average or cumulative rewards across all root-to-leaf paths in the synthesis tree [10, 21], overlooking a critical insight: a synthetic route is only valid if every root-to-leaf path terminates at a purchasable compound. Even a single unsuccessful path invalidates the entire synthetic route, making the worst-performing path the limiting factor in overall performance. This observation leads us to introduce a new “worst-path” objective that focuses explicitly on improving the most challenging path in the synthetic route.
Building on this novel formulation, we develop Interactive Retrosynthesis Planning (InterRetro) — a framework for multi-step retrosynthesis that learns to generate complete synthetic routes without search at inference time. InterRetro treats a single-step model as an agent operating within the tree MDP, recursively decomposing molecules into reactants through environment interactions. The agent is improved via weighted self-imitation of its past successful decisions to encourage shallow synthesis trees, effectively bootstrapping its performance. In particular, InterRetro operates in three key steps. Firstly, the agent interacts with the tree MDP to construct complete synthetic routes. Secondly, it identifies successful subtrees whose leaf nodes correspond to commercially available compounds. Finally, it fine-tunes the policy to imitate decisions from these subtrees, with theoretical guarantees of monotonic improvement and convergence to a unique optimal value function. This iterative self-improvement allows InterRetro to eliminate the need for real-time search while maintaining high planning quality.
@@ 2 Related Work
The fundamental task in retrosynthesis is single-step prediction, which identifies the reactants that produce a given target molecule. Early approaches are template-based, where reaction templates are extracted from patents and literature data [20], and a model selects the most appropriate template to decompose a molecule [2, 3]. Later methods propose template-free strategies, directly mapping the SMILES string of the target molecule to that of the reactants without relying on predefined transformation patterns [8, 26]. Recent studies introduce semi-template-based approaches, which combine both paradigms by first identifying intermediate structures and then completing them into full reactants — either by generating leaving groups [25], SMILES tokens [5, 32], or graph edits [30, 38]. While these single-step methods achieve high accuracy in predicting known reactions, they typically optimise for likelihood of historical reactions rather than synthesisability from commercially available compounds. In contrast, our approach explicitly favours reactions that lead to purchasable building blocks, enhancing downstream planning efficiency and practical applicability.
Multi-step Planning.
@@
Multi-step planning aims to construct a synthesis tree rooted at the target molecule, expanded by a single-step model, and terminating at commercially available compounds. Existing methods rely on heuristic search to find viable synthetic routes. For example, [6] formulates the retrosynthesis problem as an AND/OR tree search and adopts proof-number search to find the optimal solution; Retro* [1] explores the AND/OR tree using A* search and proposes value estimation methods in this setting; To improve exploration, [23] and [37] adopt MCTS, which better balances exploitation and exploration compared to A*. These search-based approaches, however, require extensive computation at inference time, often necessitating hundreds of model calls per molecule. Recent efforts such as [9] and [10] address this limitation by fine-tuning the single-step model using successful trajectories collected during the planning phase, reducing search iterations and improving synthesis success rates. Yet, these methods still ultimately rely on search during inference, merely reducing rather than eliminating this computational burden.
Our method is closely related to [21], which frames retrosynthesis as a sequential decision process and learns a policy via RL. While prior methods typically optimise the cumulative reward across the entire synthesis tree and use value-based criteria to select reactions, we take a fundamentally different approach. We introduce a novel worst-path objective that focuses on the most challenging branch of the synthesis tree, and develop a weighted imitation learning algorithm that constrains the model’s support and promotes shallower synthesis trees. This formulation enables our method to generate complete synthetic routes without any search at inference time, while maintaining high success rates.
Self-imitation.
@@
Self-imitation learning improves policy performance by encouraging the agent to replicate its own high-return past behaviours [15, 16]. A central aspect of self-imitation is the use of support constraints — the learnt policy is regularised to remain close to the data-generating policy, which stabilises training and mitigates distributional shift [16, 29]. This principle has been effectively applied in offline and offline-to-online reinforcement learning [14, 24, 27], where self-imitation facilitates safe policy improvement without extensive exploration. In our work, we apply self-imitation to retrosynthesis planning by initialising the single-step model to generate reactions from the dataset and refining it through its past successful decisions, ensuring that its proposed reactions remain chemically plausible throughout training while progressively favouring those that lead to commercially available building blocks.
@@ 3 Retrosynthesis via Worst-path Optimisation in Tree MDPs
Prior work in multi-step retrosynthesis typically minimises the total cost of a synthetic route by optimising cumulative rewards across all root-to-leaf paths in the synthesis tree. In contrast, our worst-path objective targets the weakest path, ensuring viability by requiring all paths to terminate at purchasable compounds. To efficiently optimise this criterion, we propose a weighted self-imitation algorithm that leverages successful trajectories while prioritizing improvement on failure-prone paths.
@@ 3.1 Tree-structured MDPs
We formalise retrosynthesis planning as a tree-structured Markov Decision Process (tree MDP), denoted by $\langle\mathcal{S},\mathcal{A},\mathcal{T},r,\mathcal{S}_{bb}\rangle$. Each state $s\in\mathcal{S}$ represents a molecule, each action $a\in\mathcal{A}$ represents a chemical reaction. Unlike standard MDPs where a transition yields a single successor, in retrosynthesis a reaction decomposes a molecule into multiple reactants. To capture this, we denote the set of all subsets of $\mathcal{S}$ (power set) as $2^{\mathcal{S}}$, and define a branching transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow 2^{\mathcal{S}}$, which maps a molecule $s$ and a reaction $a$ to a finite set of reactants. These reactants become the children of $s$ in the synthesis tree. The reward function $r:\mathcal{S}\rightarrow\mathbb{R}$ assigns a numerical reward to each state, and $\mathcal{S}_{bb}\subset\mathcal{S}$ denotes the set of commercially available building blocks. A synthesis tree is successful only if all of its leaf nodes belong to $\mathcal{S}_{bb}$. A policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ specifies a probability distribution over actions for each state $s$, determining how the retrosynthesis tree expands. Throughout this paper, we denote a complete retrosynthesis tree by $\tau$ and let $P(\tau)$ represent the set of all root-to-leaf paths in $\tau$.
@@
3.2 Worst-path Objective
The width of synthesis trees is naturally bounded as most molecules can be synthesised from a small number of reactants. As shown in Figure 2, $98.6\%$ of reactions in the USPTO-50k dataset involve three or fewer reactants. This observation indicates that route quality is primarily determined by depth rather than branching factor. Therefore, we evaluate a synthesis tree by its longest path from root to leaf. While our objective emphasises depth, it inherently addresses width by requiring that all reactants from a chosen reaction must lead to viable continuations. Figure 3 illustrates this concept: the left tree successfully synthesises molecule $\mathrm{A}$ with its value determined by the longest path, while the right tree fails due to an unsynthesisable intermediate, resulting in a return of 0.
Formally, we define a reward function, $r(s)=\mathds{1}\bigl{(}s\in\mathcal{S}_{bb}\bigr{)}$, which assigns a reward of $1$ if molecule $s$ is a commercially available building block, and 0 otherwise. Since $r(s_{t})$ is non-zero only at terminal states (i.e., leaf nodes), any root-to-leaf path $p=(s_{0},a_{0},s_{1},\ldots,s_{T})$ has return
@@ 3.2 Worst-path Objective
where $\gamma\in(0,1)$ is the discount factor. Thus a successful path (with $r(s_{T})=1$) receives value $\gamma^{T}$, sharply penalising deeper decompositions. Dead-end paths (where $r(s_{T})=0$) have zero return. We then define our worst-path objective as
@@
This objective captures the intuition that a synthetic route is only as viable as its weakest link: if any path leads to a dead-end, the entire tree’s value is zero. Maximising this worst-path return ensures every branch terminates at building blocks via the shortest possible routes. We also show the benefit of the worst-path objective over a best-path objective through an illustrative example in Appendix A.
Figure 3: Examples illustrating the tree MDP formulation. Each non-leaf node represents a molecule that is decomposed into one or more reactants. Left tree: A successful synthetic route for target molecule $\mathrm{A}$. It contains 44 root-to-leaf paths: $P(\tau)=\{\mathrm{ABD},\mathrm{ABEF},\mathrm{ABEGH},\mathrm{AC}\}$. Since all leaf nodes are building blocks, each path receives a value of $\gamma^{T}$ where $T$ is the path length. The tree’s overall value is $\min_{p\in P(\tau)}\{\gamma^{2},\gamma^{3},\gamma^{4},\gamma\}=\gamma^{4}$, determined by the longest path. Right tree: A failed synthesis attempt for molecule $\mathrm{A}$. One of its paths, $\mathrm{ABEG}$, terminates at $\mathrm{G}$, which is not a building block. This gives path $\mathrm{ABEG}$ a value of 0, making the tree’s overall value $\min_{p\in P(\tau)}\{\gamma^{2},\gamma^{3},0,\gamma\}=0$, illustrating why a single failing path invalidates the entire route.
3.3 Value Estimation
To optimise the worst-path objective, the agent interacts with the tree MDP and imitates its advantageous past decisions. We quantify the advantage of a decision through a Q-value function $Q^{\pi}(s,a)$: given a molecule $s$ as the root node, it estimates the expected worst-path return when first expanding the molecule with reaction $a$, and thereafter following policy $\pi$. Formally:
$Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}\left[\min_{p\in P(\tau)}\sum_{t=0}^{T}\gamma^{t}r(s_{t})\;\Big{|}\;s_{0}=s,a_{0}=a\right],$ (3)
where $s_{0}$ and $a_{0}$ denote the root node and the reaction applied to it, respectively. Similarly, the value function $V^{\pi}(s)$ represents the expected worst-path return when expanding a molecule $s$ using policy $\pi$ throughout:

### v2

@@ Abstract
Retrosynthesis planning aims to decompose target molecules into available building blocks, forming a synthetic tree where each internal node represents an intermediate compound and each leaf ideally corresponds to a purchasable reactant. However, this tree becomes invalid if any leaf node is not a valid building block, making the planning process vulnerable to the “weakest link” in the synthetic route. Existing methods often optimise for average performance across branches, failing to account for this worst-case sensitivity. In this paper, we reframe retrosynthesis as a worst-path optimisation problem within tree-structured Markov Decision Processes (MDPs). We prove that this formulation admits a unique optimal solution and provides monotonic improvement guarantees. Building on this insight, we introduce Interactive Retrosynthesis Planning (InterRetro), a method that interacts with the tree MDP, learns a value function for worst-path outcomes, and improves its policy through self-imitation, preferentially reinforcing past decisions with high estimated advantage. Empirically, InterRetro achieves state-of-the-art results – solving 100% of targets on the Retro*-190 benchmark, shortening synthetic routes by 4.9%, and achieving promising performance using only 10% of the training data.
@@ 1 Introduction
Retrosynthesis aims to identify the reactants needed to synthesise a target molecule with desired properties. As a fundamental task in computer-aided molecular design, retrosynthesis underpins progress in drug discovery and materials science dong2021deep; zhong2024recentadvances. A key challenge in retrosynthesis is single-step prediction, which involves predicting the reactants for a given product (illustrated in Figure 1). While recent approaches using supervised learning have achieved human-level accuracy in this task chen2021localretro; zhong2023graph2edits, their practical applicability is limited, as the suggested reactants are often commercially unavailable. Unlike single-step prediction, multi-step planning recursively decomposes the target into simpler intermediates, aiming to construct a synthetic route whose leaf nodes are purchasable compounds. This process naturally forms a sequential decision-making problem, where early decisions influence future steps and the overall outcome hong2023egmcts; schreck2019learning.
Most methods tackle this decision-making problem using heuristic search, with Monte Carlo Tree Search (MCTS) being widely adopted for its ability to balance exploration and exploitation roucairol2024comparing; segler2018planning. In this framework, a single-step retrosynthesis model suggests candidate reactions and predicts the most likely reactants for a given product. During simulations, MCTS selects reactions by weighing their estimated value against their exploration potential. After simulating a complete synthetic route, the values of the decisions made along the path are updated to guide future searches more effectively. Building on this foundation, recent studies have proposed various enhancements to improve the exploration–exploitation trade-off or to stabilise value estimation chen2020retrostar; hong2023egmcts; xie2022retrograph; zhao2024meea. However, these methods heavily rely on time-consuming real-time search, requiring hundreds of model calls for each molecule, limiting their practical utility in large-scale molecular design scenarios.
To improve search efficiency, recent work fine-tunes pre-trained single-step models by imitating decisions made during successful heuristic search trajectories kim2021retrostarplus; liu2023pdvn; yu2022grasp. The key insight is that reaction choices selected after look-ahead planning are often more reliable than those proposed by the original model schrittwieser2021reanalysis; wang2024goplan. By training on these improved decisions, the resulting policies can propose more plausible reactions and accelerate the search for viable synthetic routes. However, this fine-tuning process adapts the model to the distribution of molecules encountered during search, rather than those seen in direct inference. As a result, the model may perform well when used within the search loop but struggle to construct full synthetic routes independently, limiting its utility in settings where fast, search-free planning is desired.
@@
In this paper, we propose a novel perspective: reframing retrosynthesis planning as a worst-path optimisation problem in tree-structured Markov Decision Processes (tree MDPs). We observe that existing approaches typically optimise for average or cumulative rewards across all root-to-leaf paths in the synthetic tree liu2023pdvn; schreck2019learning, overlooking a critical insight: a synthetic route is only valid if every root-to-leaf path terminates at a purchasable compound. Even a single unsuccessful path invalidates the entire synthetic route, making the worst-performing path the limiting factor in overall performance. This observation leads us to introduce a new “worst-path” objective that focuses explicitly on improving the most challenging path in the synthetic route.
Building on this novel formulation, we develop Interactive Retrosynthesis Planning (InterRetro) — a framework for multi-step retrosynthesis that learns to generate complete synthetic routes without search at inference time. InterRetro treats a single-step model as an agent operating within the tree MDP, recursively decomposing molecules into reactants through environment interactions. The agent is improved via weighted self-imitation of its past successful decisions to encourage shallow synthetic trees, effectively bootstrapping its performance. In particular, InterRetro operates in three key steps. Firstly, the agent interacts with the tree MDP to construct complete synthetic routes. Secondly, it identifies successful subtrees whose leaf nodes correspond to commercially available compounds. Finally, it fine-tunes the policy to imitate decisions from these subtrees, with theoretical guarantees of monotonic improvement and convergence to a unique optimal value function. This iterative self-improvement allows InterRetro to eliminate the need for real-time search while maintaining high planning quality.
@@ 2 Related Work
Single-step Prediction. The fundamental task in retrosynthesis is single-step prediction, which identifies the reactants that produce a given target molecule. Early approaches are template-based, where reaction templates are extracted from patents and literature data schneider2016USPTO, and a model selects the most appropriate template to decompose a molecule chen2021localretro; coley2017retrosim. Later methods propose template-free strategies, directly mapping the SMILES string of the target molecule to that of the reactants without relying on predefined transformation patterns irwin2022chemformer; tu2022graph2smiles. Recent studies introduce semi-template-based approaches, which combine both paradigms by first identifying intermediate structures and then completing them into full reactants — either by generating leaving groups somnath2021graphretro, SMILES tokens han2024editretro; yan2020retroxpert, or graph edits wang2023retroexplainer; zhong2023graph2edits. While these single-step methods achieve high accuracy in predicting known reactions, they typically optimise for likelihood of historical reactions rather than synthesisability from commercially available compounds. In contrast, our approach explicitly favours reactions that lead to purchasable building blocks, enhancing downstream planning efficiency and practical applicability.
@@
Multi-step Planning. Multi-step planning aims to construct a synthetic tree rooted at the target molecule, expanded by a single-step model, and terminating at commercially available compounds. Existing methods rely on heuristic search to find viable synthetic routes. For example, heifets2021construction formulates the retrosynthesis problem as an AND/OR tree search and adopts proof-number search to find the optimal solution; Retro* chen2020retrostar explores the AND/OR tree using A* search and proposes value estimation methods in this setting; To improve exploration, segler2018planning and zhao2024meea adopt MCTS, which better balances exploitation and exploration compared to A*. These search-based approaches, however, require extensive computation at inference time, often necessitating hundreds of model calls per molecule. Recent efforts such as kim2021retrostarplus and liu2023pdvn address this limitation by fine-tuning the single-step model using successful trajectories collected during the planning phase, reducing search iterations and improving synthesis success rates. Yet, these methods still ultimately rely on search during inference, merely reducing rather than eliminating this computational burden.
@@
Self-imitation Learning. Self-imitation learning improves policy performance by encouraging the agent to replicate its own high-return past behaviours oh2018selfimitation; peng2019awr. A central aspect of self-imitation is the use of support constraints — the learned policy is regularised to remain close to the data-generating policy, which stabilises training and mitigates distributional shift peng2019awr; wang2018marwil. This principle has been effectively applied in offline and offline-to-online reinforcement learning nair2021awac; siegel2020abm; wang2025lom, where self-imitation facilitates safe policy improvement without extensive exploration. In our work, we apply self-imitation to retrosynthesis planning by initialising the single-step model to generate reactions from the dataset and refining it through its past successful decisions, ensuring that its proposed reactions remain chemically plausible throughout training while progressively favouring those that lead to commercially available building blocks.
@@ 3 Retrosynthesis via Worst-path Optimisation in Tree MDPs
Prior work in multi-step retrosynthesis typically minimises the total cost of a synthetic route by optimising cumulative rewards across all root-to-leaf paths in the synthetic tree. In contrast, our worst-path objective targets the weakest path, ensuring viability by requiring all paths to terminate at purchasable compounds. To efficiently optimise this criterion, we propose a weighted self-imitation algorithm that leverages successful trajectories while prioritising improvement on failure-prone paths.
@@ 3.1 Tree-structured MDPs
We formalise the retrosynthesis problem as a tree-structured Markov Decision Process (tree MDP), denoted by $\langle\mathcal{S},\mathcal{A},\mathcal{T},r,\mathcal{S}_{bb}\rangle$. Each state $s\in\mathcal{S}$ represents a molecule, and each action $a\in\mathcal{A}$ corresponds to a chemical reaction. Unlike standard MDPs, where each transition leads to a single successor state, a chemical reaction may decompose a molecule into multiple reactants. To capture this branching structure, we define the power set of all possible molecules as $2^{\mathcal{S}}$ and introduce a branching transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow 2^{\mathcal{S}}$, which maps a product molecule $s$ and a reaction $a$ to a set of reactants. These reactants become the children of $s$ in the synthetic tree. The reward function $r:\mathcal{S}\rightarrow\mathbb{R}$ assigns a numerical value to each molecule, and $\mathcal{S}_{bb}\subset\mathcal{S}$ denotes the set of commercially available building blocks. A synthetic tree is considered successful only if all of its leaf nodes belong to $\mathcal{S}_{bb}$. A policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ defines a probability distribution over feasible reactions for a molecule $s$, thereby determining how the synthetic tree expands. Throughout this paper, we denote a complete synthetic tree by $\tau$, and let $P(\tau)$ represent the set of all root-to-leaf paths within $\tau$. The tree MDP formulation is illustrated in Figure 3.
3.2 Worst-path Objective
@@
The width of synthetic trees is naturally bounded, as most molecules can be synthesised from only a few reactants. As shown in Figure 2, $98.6\%$ of reactions in the USPTO-50k dataset involve three or fewer reactants. This suggests that the quality of a synthetic route is primarily determined by its depth rather than its branching factor. Accordingly, we evaluate each synthetic tree by the length of its longest root-to-leaf path. Figure 3 illustrates this concept: the left tree successfully synthesises molecule $\mathrm{A}$, with its value determined by the longest path, whereas the right tree fails due to an unsynthesisable intermediate $G$, yielding a return of 0.
Formally, we define the reward function as
$r(s)=\begin{cases}1,&\text{if }s\in\mathcal{S}_{bb},\\ 0,&\text{otherwise}.\end{cases}$ (1)
This assigns a reward of $1$ to commercially available building blocks and 0 otherwise. Since the reward is non-zero only at terminal states (i.e., leaf nodes), the return of any root-to-leaf path $p=(s_{0},a_{0},s_{1},\ldots,s_{T})$ is
@@
where $\gamma\in(0,1)$ is the discount factor. Hence, a successful path with $r(s_{T})=1$ receives value $\gamma^{T}$, penalising deeper decompositions, while dead-end paths with $r(s_{T})=0$ yield zero return. We then define the worst-path objective as
@@
This objective captures the intuition that a synthetic route is only as viable as its weakest link: if any path leads to a dead-end, the entire tree’s value becomes zero. Maximising the worst-path return therefore encourages the policy to ensure that every branch terminates at building blocks through the shortest possible routes.
Figure 3: Examples illustrating the tree MDP formulation. Each non-leaf node represents a molecule that is decomposed into one or more reactants. Left tree: A successful synthetic route for target molecule $\mathrm{A}$. It contains 44 root-to-leaf paths: $P(\tau)=\{\mathrm{ABD},\mathrm{ABEF},\mathrm{ABEGH},\mathrm{AC}\}$. Since all leaf nodes are building blocks, each path receives a value of $\gamma^{T}$, where $T$ is the path length. The tree’s overall value is $\min_{p\in P(\tau)}\{\gamma^{2},\gamma^{3},\gamma^{4},\gamma\}=\gamma^{4}$, determined by the longest path. Right tree: A failed synthesis attempt for molecule $\mathrm{A}$. One of its paths, $\mathrm{ABEG}$, terminates at $\mathrm{G}$, which is not a building block. This gives path $\mathrm{ABEG}$ a value of 0, making the tree’s overall value $\min_{p\in P(\tau)}\{\gamma^{2},\gamma^{3},0,\gamma\}=0$, illustrating why a single failing path invalidates the entire route.
3.3 Advantage Estimation
To optimise the worst-path objective, the agent interacts with the tree-structured MDP and learns to reproduce its past advantageous decisions. In this section, we formally define the advantage to quantify how beneficial a reaction is relative to the policy’s average behaviour. We then propose to estimate this advantage using a value function learned from interaction experiences. Finally, we analyse its theoretical properties and establish the existence of an optimal policy that maximises the worst-path objective.
Given a molecule $s$ as the root node, a Q-function estimates the expected worst-path return when first expanding $s$ with reaction $a$ and subsequently following policy $\pi$:
$Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}\!\left[\min_{p\in P(\tau)}\sum_{t=0}^{T}\gamma^{t}r(s_{t})\;\Big|\;s_{0}=s,a_{0}=a\right].$ (4)
Here, $s_{0}$ and $a_{0}$ denote the root molecule and the initial reaction, respectively. Similarly, the value function $V^{\pi}(s)$ represents the expected worst-path return when all subsequent reactions follow policy $\pi$: